OpenStudy (hdrager):

Can somebody explain to me how to do this please? cos(arctan 3/4)

OpenStudy (anonymous):

gladly! it's saying what angles have a tan=3/4 ? tan is opp/adj opposite=3 adjacent=4 that means hypotenuse=5 cos is adjacent/hypotenuse knowing both adjacent and hypotenuse, cos of the angle that has a tan=3/4 would be cos(4/5) which is equal to what?

jimthompson5910 (jim_thompson5910):

@ijakez you're on the right track, but you made an error when you got to `cos(4/5)`

OpenStudy (anonymous):

the angle theta that has a tangent tan(theta)=3/4 that angle would have a cos(theta)=4/5 arccos(4/5)=? oopsies!

OpenStudy (mathmale):

The following wording might make more sense to you and others: Instead of "knowing both adjacent and hypotenuse, cos of the angle that has a tan=3/4 would be cos(4/5) which is equal to what?" "The cosine of the angle that has a tangent of 3/4 would be adj/hyp (as usual), which in this case comes out to 4/5." (Not cos (4/5). You could also do this: "Let x be the angle in question. Then tan x = 3/4 and cos x = 4/5."

OpenStudy (anonymous):

yeah my mistake